Sunland Company has a net profit margin of 7.5 percent, debt ratio of 42 percent, total assets of $4,588,700, and sales of $5,520,600. If the company has a dividend payout ratio of 74 percent, what is its sustainable growth rate? (Round answer to 1 decimal place, e.g. 17.5%.)

Sustainable growth rate enter the sustainable growth rate in percentages rounded to 1 decimal place %
Answer #1

ROE = Net income / equity

Net income = Total sales * Net profit margin = $5,520,600 * 7.5% = $414,045

Debt ratio = 0.42

Debt / total assets = 0.42 = Debt / $4,588,700

Debt = $1,927,254.

Equity = Total Assets - Total Debt

Equity = $4,588,700 - $1,927,254.

Equity = $2,661,446

ROE = $414,045 / $2,661,446

ROE = 0.1556

Retention ratio (b) = 1 - dividend payout ratio = 1 - 0.74

b = 0.26

Sustainable growth rate = (ROE * b) / [1 - (ROE*b)]

= 0.0404486 / (0.95955)

= 0.042

or 4.2%